Help one more time or maybe 2 at the most. I hate word problems!!
A photographer is mixing a 5% acetic acid solution with a 10% acetic solution to get 5 liters of a 7%solution. How many liters of the 5 % and 10%solutions should she add? Show work
HELP!!

Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

A photographer is mixing a 5% acetic acid solution with a 10% acetic solution to get 5 liters of a 7%solution.
How many liters of the 5 % and 10%solutions should he add?
:
let x = amt of 10% solution
the results is 5 liters, therefore
(5-x) = amt of 5% solution
:
.10x + .05(5-x) = .07(5)
.10x + .25 - .05x = .35
.10x - .05x = .35 - .25
.05x = .10
x = .10/.05
x = 2 liters of the 10% solution
then obviously
5 - 2 = 3 liters of the 5% solution

Answer by ikleyn(52790)

.
You hate word problems because you are not trained to solve them.

Meanwhile, there is a lot of lessons on almost all types of word problems in this site, prepared specially for you.

They are in the online textbook
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.
which is available for you for free!

Yes, free of charge.

There is the entire section "Word problems" in the textbook with detailed solutions.

Read them and become an expert in this area.

You will love them!